OpenStudy (anonymous):

What is the factored form of 16x2 – 16x – 12?

OpenStudy (tanner23456):

4(4x^2-4x-3) 4x^2-4x-3 first *last = 4*3 12 factor out 4*3 2*2*3 add to get 4 2*3 - 2 6-2 = 4 4x^2 -6x+2x-3 (4x^2-6x)(2x-3) 2x(2x-3)+1(2x-3) (2x+1)(2x-3)4

OpenStudy (anonymous):

thanks